[figure: 9.7 cm]

Specific features in electrical- and magnetic resistivity of RFe_{11}Ti single crystals and their hydrides

I.S.Tereshina^{a,b}, S.A.Nikitin^{a,c}, G.A.Bezkorovajnaya^{a}, W.Suski^{b,c}, D.Badurski^{b} aDepartment of Physics, Moscow State University, 119992 Moscow, Russia

bPolish Academy of Sciences, W. Trzebiatowski Institute of Low Temperature and

Structure Research, P.O. Box 1410, 50-950 Wroclaw 2, Poland cInternational Laboratory of High Magnetic Fields and Low Temperatures, 95 Gajowicka str., 53-421 Wroclaw, Poland

The temperature dependences of the electrical resistivity (ρ) of RFe_{11}TiH_{x} (R = Y, Dy, Er; x = 0, 1) has been investigated. They exhibit metallic character. Hydrogenation leads to decrease of electrical resistivity. Magnetic resistivity and its anomalies in the vicinity of spin-reorientation transitions (SRT) in single crystals of RFe_{11}Ti (R = Dy and Er) are investigated in the temperature range 4.2-130 K and in magnetic fields up to 8 T. The ErFe_{11}Ti shows an SRT at T = 50 K from the c axis to a non-collinear magnetic structure at low temperatures which result in sharp increase of ρ(H) at low magnetic fields. Also, anomalous increases of magnetization (first-order magnetization process) and ρ(T) are obtained for DyFe_{11}Ti when a field of about 1-2 T is applied along [001] axis at a temperature in the range 4.2-130 K.

[figure: 13.4 cm]

Subject category :

5. Phase Transitions and Critical Phenomena

Presentation mode : poster

Corresponding author : S.A.Nikitin

Address for correspondence : M.V.Lomonosov Moscow State University, Department of Physics, KOFEF Vorobyevy Gory, 119992 Moscow, Russia

Email address : nikitin@rem.phys.msu.su